the length of sallys garden is 4 meters greater than 3 times the width. the perimeter of her garden is 72 meters. what is the length of her garden?
L=4+3(w)?????

Answer by Sarpi(32)


The garden has a rectangular shape. Therefore the perimeter (P) = 2(L+W); where P= 72, L = 4 + 3W and W is not known.
P = 2(L+W)
P = 2L+ 2W
72 = 2(4+3W) + 2W
72 = 8 + 6W + 2W
72 = 8 + 8W
72 - 8 = 8W
64 = 8W
W = 64/8
W = 8 meters.
Therefore, the length = 4 + 3*8 = 28 meters.
Verification:
P = 2(L+W)
p = 2(28+8)
P = 72 meters.